OpenStudy (anonymous):

f(x)=(6e^(x)−4)/(14e^(x)+6). Find the inverse function and its domain (interval notation)

OpenStudy (mathmale):

By f(x)=(6e^(x)−4)/(14e^(x)+6), do you mean\[y=\frac{ 6e^x-4 }{ 14e^6+6 }?\]

OpenStudy (anonymous):

oh it's this: \[f(x)=\frac{ (6e^x-4) }{ (14e^x +6) }\]

OpenStudy (anonymous):

@mathmale

OpenStudy (mathmale):

thanks for verifying the problem statement. Writing your version as a fraction, as I did, eliminates the need for parentheses.\[y=\frac{ 6e^x-4 }{ 14e^x+6 }\]

OpenStudy (mathmale):

You're correct; the problem was given to you labeled f(x). The next step is to replace the "f(x)=" label with "y=" ... which I have done. Next, interchange x and y: where you see x, write y; where you see y, write x instead. Pls do that now.

OpenStudy (anonymous):

\[x=\frac{ 6e^y -4}{ 14e^y+6 }?\]

OpenStudy (mathmale):

Yes. Now, please solve this new equation for e^y.

OpenStudy (anonymous):

I'm not sure how to do this part... Do I need to change e to ln?? But I know that I can move the denominator to the right hand side: \[x(14e^y+6)=6e^y-4\]

OpenStudy (mathmale):

You're on the right track! However, you still need to solve this equation for e^y. Mult. the quantity within parentheses by x. Next, group the e^y terms on the left side and the remaining terms on the right side. Solve for e^y.

OpenStudy (anonymous):

So like this: \[6x+4=6e^y-14xe^y\] \[6x+4=e^y(6-14x)\]

OpenStudy (mathmale):

Looks good, and I appreciate your showing your work as you have. Please divide both sides of this equation by (6-14x) to isolate e^y. Note that the left side can be reduced. How?

OpenStudy (anonymous):

It will like this: \[\frac{ 6x+4 }{ 6-14x }=e^y\] \[\frac{(3x+2) }{(3-7x) }=e^y\]

OpenStudy (mathmale):

Summary: replace f(x) with y. 2. Interchange x and y 3 Solve for y 4. replace y with \[f ^{-1}(x)=\]

OpenStudy (mathmale):

We overlooked one step. \[\frac{(3x+2) }{(3-7x) }=e^y\]

OpenStudy (mathmale):

... is fine, but we need to go all the way in solving for y. Apply the ln operator to both sides of this most recent question. That will isolate y on the right side.

OpenStudy (anonymous):

oohh... \[\ln \frac{ 3x+2 }{ 3-7x }=(e^y)\ln\]

OpenStudy (anonymous):

\[y=\ln(\frac{ 3x+2 }{ 3-7x })\]

OpenStudy (mathmale):

Looks good! You may either leave it that way, or write:\[y=\ln (3x+2)-\ln (3-7x)\]

OpenStudy (anonymous):

Oh I see... I forgot the properties for "e" and "ln"

OpenStudy (mathmale):

You could check by either: 1) substitute this result (right side only) into the original equation, replacing x in the original equation, or 2) substitute the original equation (right side only) into our most recent "answer." If the end result is merely "x"

OpenStudy (mathmale):

...then you have found the correct inverse function of the original f(x).

OpenStudy (anonymous):

Oh okay... How would we find the domain (interval notation)??

OpenStudy (mathmale):

"ln x" shows up 2x in your latest result.

OpenStudy (mathmale):

Recall that the domain of ln x is (0, +inf). Example: Set 3x+2>0 and solve for x. Do the same for the second argument: 3-7x>0. graph both sets. The overall, final domain is the set of all x where the 2 previous sets overlap one another.

OpenStudy (anonymous):

So the domain would be: \[x >-2/3\] and \[x >3/7\]

OpenStudy (mathmale):



OpenStudy (mathmale):

You must now graph your 2 results. Determine the interval on which these two results overlap. ONLY the overlapping part constitutes the domain for your original inequality.

OpenStudy (mathmale):

Glad to have the chance to work with you. I must get off the 'Net now, but if you have further questions about this probelm, just leave a note for me; I will see it in the morning. Good night!

OpenStudy (anonymous):

I think I got it. Thank you so much! I learned a lot! :)

OpenStudy (anonymous):

But I'm not sure what you mean by overlapping... I think the interval notation is (-2/3, 3/7) @mathmale